Displaying report 1-1 of 1.
Reports until 17:59, Monday 29 July 2024
GQuEST Output Filter Cavities
Torrey Cullen - posted 17:59, Monday 29 July 2024 (11788)
OFC2 Aligned

[Ian,Torrey]

OF2 is aligned! Process is the same as described in the past. More difficult having the input coupler as the static piezo mirror this time. Attached is a cavity scan for both cavities. It seems that the new one is much better aligned, although we should double check on a faster TRANS PD.

We went to lock it and noticed that the tank circuit on the 775 EOM had a wire broken. We resoldered in a new metal box to prevent this in the future.

Tomorrow we will take back the fourth moku and outfit it to control the second filter cavity as well as align the 1550 test light.

Images attached to this report
Displaying report 1-1 of 1.